# Robust assessment of asymmetric division in colon cancer cells

### v2

@@ IV.1 Cell Culture
CCD-18Co, colon fibroblast cell line, were purchased from ATCC (Manassas, VA, USA, CRL-1459) and maintained in complete culture media DMEM containing $10\%$ FBS, penicillin/streptomycin plus glutamin, nonessential amino acids (NEAA) and sodium pyruvate, all $1/100$ dilution.
All cells were kept in culture at $37^{\circ}C$ in $5\%$ CO2 and passaged according to the experimental protocol. For all experiments cells were harvested, counted, and washed twice in serum-free solutions and resuspended in room temperature (RT) PBS w/o salts for further staining.
In detail, to detach cells from culture flasks we used Trypsin EDTA $0.25\%$. Culture media was removed and the flasks were washed once with PBS, then Trypsin EDTA was added and allowed to work for $1$ min in an incubator at $37^{\circ}C$. Once the cells appeared detached the flasks were mechanically agitated to facilitate cells’ loss of adhesion. Trypsin was then inactivated with one volumes of complete media and cells were collected, pelleted, and washed a second time in PBS. To determine cell viability, prior to dye staining, the collected cells were counted with the hemocytometer using Trypan Blue, an impermeable dye not taken up by viable cells.
@@ IV.2 Flow Cytometry and Cell Sorting
To track cell proliferation by dye dilution for establishing the progeny of a mother cell, cells were stained with CellTrace VioletTM (CTV). The CTV dye staining (C34557, Life Technologies, Paisley, UK), used to monitor multiple cell generations, was performed according to the manufacturer’s instruction, diluting the CTV 1/1000 in 0.5-1 ml of PBS for 20 min in a water bath at $37^{\circ}C$, mixing every 10 min. Afterward, 5x complete media was added to the cell suspension for an additional 5 min incubation before the final washing in PBS.
Labeled cells were sorted using a FACSAriaIII (Becton Dickinson, BD Biosciences, USA) equipped with Near UV 375nm, 488nm, 561nm, and 633nm lasers and FACSDiva software (BD Biosciences version 6.1.3). Data were analyzed using FlowJo software (Tree Star, version 10.7.1). Briefly, cells were first gated on single cells, by doublets exclusion with morphology parameters area versus width (A versus W), both side and forward scatter. The unstained sample was used to set the background fluorescence. The sorting gate was set around the max peak of fluorescence of the dye distribution. In this way, the collected cells were enriched for the highest fluorescence intensity. Following isolation, an aliquot of the sorted cells was analyzed with the same instrument to determine the post-sorting purity and population width, resulting in an enrichment $\rangle$ 97 % for each sample.
To monitor multiple cell division, the sorted cell population was seeded in 12 wells plates (Corning, Kennebunk, ME, USA) at cell density bewteen $[30-70]\cdot 10^{3}$ cells/well, according to the experiment, and kept in culture for up to 84 h. Each well corresponds to a time point of the acquisition and cells in culture were analyzed every 24, 36, 48, 60, 72, 84 h by the LSRFortessa flow cytometer. In order to set the time zero of the kinetic, prior culturing, a tiny aliquot of the collected cells was analyzed immediately after sorting at the flow cytometer. The unstained sample was used to set the background fluorescence as described above.
@@ IV.3 Time-lapse Microscopy
To better investigate the cell proliferation dynamics, we performed time-lapse experiments, for up to 3 days. In our previous paper peruzzi2021asymmetric we verified that cell growth is not affected by different dyes combination. Therefore, in the present work, we used both CellTrace YellowTM (C34573 A, Life Technologies, Paisley, UK) and CellTrace Far RedTM (C34572, Life Technologies, Paisley, UK), to stain the cytoplasm for microscopy analysis.
[…]
@@
M.M. and G.G. conceived research; D.C., G.P., and S.S. performed cell biology experiment and flow cytometry measurements. D.C., C.G., V.d.T., and M.M. performed timelapse microscopy experiments; F.G. and G.R. contributed additional ideas and helped interpret results; D.C. analyzed data; M.M. performed analytical calculations; D.C. and G.G. performed numerical simulations and statistical analysis; all authors analyzed results, and wrote and revised the paper.
@@ Acknowledgements
This research was partially funded by grants from ERC-2019-Synergy Grant (ASTRA, n. 855923); EIC-2022-PathfinderOpen (ivBM-4PAP, n. 101098989); Project ‘National Center for Gene Therapy and Drugs based on RNA Technology’ (CN00000041) financed by NextGeneration EU PNRR MUR—M4C2—Action 1.4—Call ‘Potenziamento strutture di ricerca e creazione di campioni nazionali di R&S’ (CUP J33C22001130001).

### v3

@@ IV.1 Cell Culture
CCD-18Co, colon fibroblast cell line, was purchased from ATCC (Manassas, VA, USA, CRL-1459) and maintained in complete culture media DMEM containing $10\%$ FBS, penicillin/streptomycin plus glutamin, nonessential amino acids (NEAA), and sodium pyruvate, all $1/100$ dilution.
All cells were kept in culture at $37^{\circ}C$ in $5\%$ CO2 and passaged according to the experimental protocol. For all experiments, cells were harvested, counted, and washed twice in serum-free solutions and resuspended in room temperature (RT) PBS w/o salts for further staining.
In detail, to detach cells from culture flasks, we used Trypsin EDTA $0.25\%$. Culture media was removed and the flasks were washed once with PBS, then Trypsin EDTA was added and allowed to work for $1$ min in an incubator at $37^{\circ}C$. Once the cells appeared detached, the flasks were mechanically agitated to facilitate the cells’ loss of adhesion. Trypsin was then inactivated with one volume of complete media, and cells were collected, pelleted, and washed a second time in PBS. To determine cell viability, prior to dye staining, the collected cells were counted with the hemocytometer using Trypan Blue, an impermeable dye not taken up by viable cells.
@@ IV.2 Flow Cytometry and Cell Sorting
To track cell proliferation by dye dilution for establishing the progeny of a mother cell, cells were stained with CellTrace VioletTM (CTV). The CTV dye staining (C34557, Life Technologies, Paisley, UK), used to monitor multiple cell generations, was performed according to the manufacturer’s instructions, diluting the CTV 1/1000 in 0.5-1 ml of PBS for 20 min in a water bath at $37^{\circ}C$, mixing every 10 min. Afterward, 5x complete media was added to the cell suspension for an additional 5 min incubation before the final washing in PBS.
Labeled cells were sorted using a FACSAriaIII (Becton Dickinson, BD Biosciences, USA) equipped with Near UV 375nm, 488nm, 561nm, and 633nm lasers and FACSDiva software (BD Biosciences version 6.1.3). Data were analyzed using FlowJo software (Tree Star, version 10.7.1). Briefly, cells were first gated on single cells, by doublet exclusion with morphology parameters area versus width (A versus W), both side and forward scatter. The unstained sample was used to set the background fluorescence. The sorting gate was set around the maximum peak of fluorescence of the dye distribution. In this way, the collected cells were enriched for the highest fluorescence intensity. Following isolation, an aliquot of the sorted cells was analyzed with the same instrument to determine the post-sorting purity and population width, resulting in an enrichment of $\rangle$ 97 % for each sample.
To monitor multiple cell divisions, the sorted cell population was seeded in 12-well plates (Corning, Kennebunk, ME, USA) at a cell density between $30-70]\cdot 10^{3}$ cells/well, according to the experiment, and kept in culture for up to 84 h. Each well corresponds to a time point of the acquisition, and cells in culture were analyzed every 24, 36, 48, 60, 72, and 84 h by the LSRFortessa flow cytometer. In order to set the time zero of the kinetic, prior to culturing, a tiny aliquot of the collected cells was analyzed immediately after sorting at the flow cytometer. The unstained sample was used to set the background fluorescence as described above.
@@ IV.3 Time-lapse Microscopy
To better investigate the cell proliferation dynamics, we performed time-lapse experiments for up to 3 days. In our previous paper peruzzi2021asymmetric , we verified that cell growth is not affected by different dye combinations. Therefore, in the present work, we used both CellTrace YellowTM (C34573 A, Life Technologies, Paisley, UK) and CellTrace Far RedTM (C34572, Life Technologies, Paisley, UK) to stain the cytoplasm for microscopy analysis.
Low passage cells at around 60 $\%$ confluency were counted and stained with CellTrace dyes 1/500 and plated on IBIDI cell imaging chambers ($\mu$-Slide 4 and 8 wells) at a low cell density of $10^{3}$ cells/well. After overnight incubation, the chamber is transferred to the inverted microscope adapted with an incubator to keep cells in appropriate growing conditions. Brightfield and fluorescent confocal image stacks were acquired with a 20x air objective (Olympus, Shinjuku, Japan) and Zen Microscopy Software (Zeiss, Oberkochen, Germany), every 20 min.
@@
M.M. and G.G. conceived the research; D.C., G.P., and S.S. performed the cell biology experiment and flow cytometry measurements. D.C., C.G., V.d.T., and M.M. performed timelapse microscopy experiments; F.G. and G.R. contributed additional ideas and helped interpret results; D.C. analyzed data; M.M. performed analytical calculations; D.C. and G.G. performed numerical simulations and statistical analysis; all authors analyzed results and wrote and revised the paper.
@@ Acknowledgements
This research was partially funded by grants from ERC-2019-Synergy Grant (ASTRA, n. 855923); EIC-2022-PathfinderOpen (ivBM-4PAP, n. 101098989); Project ‘National Center for Gene Therapy and Drugs based on RNA Technology’ (CN00000041) financed by NextGeneration EU PNRR MUR—M4C2—Action 1.4—Call ‘Potenziamento strutture di ricerca e creazione di campioni nazionali di R&S’ (CUP J33C22001130001). Project FIS2023-02957 (CUP B53C24009530001), ’Fathoming oUt the role of partitioninG noIse in cancer epithelial-mesenchymal Transitions (FUGIT), funded with the contribution of the Italian Ministry of University and Research pursuant to Ministerial Decree No. 1236 of August 1, 2023 - FIS 2 CALL.